OpenStudy (anonymous):

Give an example of a rational function that has a horizontal asymptote of y = 2/9.

OpenStudy (tkhunny):

Not even going to try? Three Examples \(y = \dfrac{2}{9}\) \(y = \dfrac{2(x-1)}{9(x+1)}\) \(y = \dfrac{2(x^{2}-1)}{9(x+1)(x+3)}\) With only one constraint, you should be able to create these as fast as you can write.